OpenStudy (anonymous):

What is the partial sum 25 ∑ -5+13i i=12

OpenStudy (anonymous):

3,297 3,630 3,976 4,335

OpenStudy (amistre64):

the sum notation is a linear transformation ... if i recall the definition of a linear transformation correctly

OpenStudy (amistre64):

∑ -5+13i -5∑ + 13 ∑ i -5(range) + 13 (range(n*(n+1)/2))

OpenStudy (amistre64):

range would be defined as: high - low + 1

OpenStudy (amistre64):

or, we could translate i from 12 to 0 and add 12 to each i to keep it the same

OpenStudy (amistre64):

25 ∑ -5+13i i=12 13 ∑ -5+13(i+12) i=0 there are so many ways we can approach this

OpenStudy (anonymous):

i still dont understand it really

OpenStudy (amistre64):

if you wanna brute math it into submission, the most fundamental means would be to copy the equation itself a few times such that for each copy you replace the "i" by the values from 12 to 25; then add them all up

OpenStudy (amistre64):

-5+13(12)= -5+13(13)= -5+13(14)= -5+13(15)= .... -5+13(25)=

OpenStudy (anonymous):

so after i get each total, do i add all the totals up?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

that is the long way to do it, but it works in the end :)

OpenStudy (anonymous):

thank you....could i do that for every equation ?

OpenStudy (amistre64):

1: -5+13(12) 2: -5+13(13) 3: -5+13(14) 4: -5+13(15) 5: -5+13(16) 6: -5+13(17) 7: -5+13(18) 8: -5+13(19) 9: -5+13(20) 10:-5+13(21) 11:-5+13(22) 12:-5+13(23) 13:-5+13(24) 14:-5+13(25) -------------- -5(14) + 13(14)(12+25)/2

OpenStudy (amistre64):

you can do it for reasonably small equations by hand. the more terms there are the better it is to know the shorter methods

OpenStudy (anonymous):

thank you. I found the answer

OpenStudy (amistre64):

cool :) what was it ?

OpenStudy (anonymous):

3297

OpenStudy (amistre64):

yay!! i got that too, so either we both did it right, or we both got it terribly wrong lol

OpenStudy (anonymous):

i have another question, and if you dont mind can you please help me?

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

go ahead and ask

OpenStudy (anonymous):

Given the sequence 3, 9, 15, which of the following is equivalent to 8 ∑ An n=5

OpenStudy (amistre64):

does it say if the sequence of arith or geoM?

OpenStudy (anonymous):

no it doesn't

OpenStudy (amistre64):

a1 = 3 a2 = 3 + 6 = 9 a3 = 9 + 6 = 15 so this looks to be that we are adding the same value each time to get the next term. we call that an arithmetic sequence

OpenStudy (amistre64):

if we can form the general equation that defines this sequence, we can then do the same thing as the we did on the first question

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

a1 = 3 a2 = 3 + 6 = 9 a3 = 3 + 6 + 6 3 + 6(2) = 15 an = 3 + 6(n-1) is the equation we seek

OpenStudy (anonymous):

okay...im going to try and solve it and see what i get

OpenStudy (amistre64):

\[\large \sum_{n=5}^{8}3+6(n-1)\]

OpenStudy (amistre64):

1: 3+6(4) 2: 3+6(5) 3: 3+6(6) 4: 3+6(7) ------- 3(4)+6(4)(4+5+6+7)

OpenStudy (anonymous):

i got 144

OpenStudy (amistre64):

12+24(22) = 12+ 528 is going to be a little bit bigger than 144

OpenStudy (anonymous):

144 is one of my answer choices . the highest one is 168

OpenStudy (amistre64):

really? 1: 3 2: 9 3: 15 4: 21 .......... 5: 27 6: 33 7: 39 8: 45 ------ 144 hmmm, i wonder what i was thinking about then :)

OpenStudy (anonymous):

lol...yea i am about to submit the answers after i finish one more question and i will let you know if it was right or not

OpenStudy (anonymous):

They were right...thank you